OpenStudy (anonymous):

Identify the type of Conic Section that has the equation 16x^2+4y^2=16 and identify its domain and range.

OpenStudy (anonymous):

I already know that it is an Ellipse. but I want to know step by step how to get the range and domain of the equation.

OpenStudy (anonymous):

@bubbles2000

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

i wouldn't know sorry :(

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@tkhunny

OpenStudy (imstuck):

It has to equal 1, first of all, so divide everything by 16 to get x^2/1 + y^2/4 = 1. Since in an ellipse, the a is always greater than the b, and since the 4 is under the y term, that is your major axis. The formula is x^2/a^2 + y^2/b^2 = 1, or x^2/b^2 + y^2/a^2 = 1. Like I said, since the 4 is under the y term, the y is the major axis with the points located at (0, 2) and (0, -2). That's the range (range is y). The minor axis, the x axis has its vertices at (1, 0) and (-1, 0). Since there are no constants in the numerators with the x^2 and y^2, this is centered at the origin. The domain, btw, is [-1,1]. Range is [2,-2]. I hope this helps!

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

@IMstuck can u help me with another question?

OpenStudy (anonymous):

The reflecting dish of a parabolic microphone has a cross-section in the shape of a parabola. The microphone itself is placed on the focus of the parabola. If the parabola is 24 inches wide and 4 inches deep, how far from the vertex should the microphone be placed?

OpenStudy (imstuck):

Let me think on think on this one...

OpenStudy (tkhunny):

@Miranda122 Post a new thread. Set it up so that: Vertex: (0,0) Point: (-12,4) Point: (12,4) Show your work on the other thread when you create it.